Extreme hardship. Kim starts to walk 3 mi to school at
7:30 A.M. with a temperature of 0°F. Her brother Bryan
starts at 7:40 A.M. on his bicycle, traveling 10 mph faster
than Kim. If they get to school at the same time, then how
fast is each one traveling?

Answer by stanbon(75887)

Kim starts to walk 3 mi to school at
7:30 A.M. with a temperature of 0°F. Her brother Bryan
starts at 7:40 A.M. on his bicycle, traveling 10 mph faster
than Kim. If they get to school at the same time, then how
fast is each one traveling?
---
Kim DATA:
distance = 3 mi ; rate = x mph ; time = 3/x hrs
==============
Bryan DATA:
distance = 3 mi ; rate = x+10 mph ; tim = 3/(x+10)
-------------------
Equation:
Kim time - Bryan time = 1/6 hr
3/x - 3/(x+10) = 1/6
----
Multiply thru by 6x(x+10) to get:
18(x+10) - 18x = x(x+10)
180 = x^2+10x
------
x^2+10x-180 = 0
-------
Positive solution:
x = [-10+sqrt(100-4*-180]/2
x = 9.32 mph (Kim speed)
---
x+10 = 19.32 mph (Bryan speed)
===================================
Cheers,
Stan H.
==========